OpenStudy (anonymous):

In one hour of the decay rate of a small amount of radioactive isotope drops from 16000 decays per second to 200 decays per second. What is the half life. Help Please!!!!

OpenStudy (anonymous):

\(R=R_0 (\frac{1}{2})^{\frac{t}{t_{1/2}}}\) this is the equation for rate change where t=time lapse, R_0 the original rate and R the rate , t1/2 =halflife sub everything and get t_1/2